OpenStudy (anonymous):

cramer's Rule 3 x3 ' -1 1 1 ' ' -9 5 3 ' ' 5 -2 -1 ' ' ' the key is -2, but my answer is 0,thank you in advance

OpenStudy (anonymous):

The answer is -2. The equation to solve is: -1(-5+6)-1(9-15)+1(18-25)

OpenStudy (anonymous):

[5 3] [-9 3] [-9 5] + (-1) [-2 -1] - (1)[5 -1] +(1)[5 -2] =-1(-5+6)-1(9-15)+1(18-25) = -1(1) -1(-6) +1(-7) =-2 ans